Fermi National Laboratory


Accelerator Update

Friday May 14
The day shift began with Operations monitoring store 3491, with the Antiproton Source (Pbar) stacking, with Recycler experts conducting studies, and with MiniBooNE and the Meson MCenter experiments taking beam.

At 9:37 AM, a Pbar expert began a stacking study to be followed by a reverse proton study.

Operations began having gate bias trouble with the Main Injector (MI) RF stations at 10:27 AM.

Around noon Operations began having beam problems in MI. An expert investigated and discovered a ramp dip. Operators then discovered and corrected a Time Line Generator (TLG) conflict.

The Recycler expert completed his study at 12:34 PM.

A Booster power supply (ORBMP) tripped off at 1:32 PM. Experts investigated and found a major problem. Operators made the tunnel ready for access at 2:20 PM.

MI suffered a vacuum burst at 3 PM. It reset okay.

Booster technicians began replacing a power amplifier for a RF station (BRF16) at 3:22 PM.

Techs completed their work in Booster and Operations established beam through Booster to the Accumulator. Pbar resumed stacking at 9:02 PM.

Operations established beam to MiniBooNE at 9:14 PM.

Saturday May 15
The midnight shift began with Operations monitoring store 3491, with Pbar stacking, and with beam to MiniBooNE and MCenter.

At 12:55 AM, Operations discovered that the beam wasn't getting to MCenter, but due to a problem with the QXR was being sent to the MI dump, which was causing MI beam problems. Operations reset the QXR and everything returned to normal operations.

Operations terminated store 3491 at 8 AM and then put the Tevatron (TeV) into shot setup.

At 8:33 AM, the Collider Detector at Fermilab (CDF) asked for an urgent access to repair a power supply. Operations stopped the shot setup. CDF finished their work at 10:32 AM.

At 10:36 AM, a TeV kicker located at sector E4 alarmed on a power supply failure. Operators called in an expert and together they topped off some cooling tanks, which reset the alarm.

Operations put the TeV into shot setup at 12:43 PM.

Operations established store 3494 at 2:22 PM with an initial luminosity of 51.52E30. Pbar resumed stacking.

The TeV suffered a multi-house quench at 2:46 PM. Experts began their investigations.

The Meson MTest experiment requested beam 3:18 PM.

At 4:11 PM, the Booster ORBMP power supply stopped putting out current. Two experts came in, repaired the supply and had it back on line by 6:11 PM. Stacking resumed about ten minutes later.

Operations reestablished beam to MCenter at 6:28 PM, but MTest quit for the day.

Operations put the TeV through an extended dry squeeze at 7:05 PM.

Operations put the TeV through a wet squeeze at 10:27 PM.

A TeV expert began a study at 10:52 PM.

Operations reported at 11:48 PM that no cause was found for the quench.

Sunday May 16
The midnight shift began with a TeV expert conducting studies, with Pbar stacking, and with MiniBooNE and MCenter taking beam.

At 1:57 AM, Operations put the TeV through another wet squeeze.

Operations put the TeV into shot setup at 2:22 AM.

At 4:25 AM, Operations established store 3500 with an initial luminosity of 52.82E30. Pbar resumed stacking. Beam problems appeared almost immediately and operators began tuning. Operators called an expert at 5:36 AM who helped with collimator adjustments.

Operations resumed sending beam to MCenter at 4:49 AM.

Operations reported at 7:42 AM that Pbar stopped accumulating antiprotons. Operators investigated and discovered that the lithium lens target shutter was closed and stuck. They called an expert who suggested that operators power cycler the motor control box. This resolved the problem at 8:20 AM and stacking resumed.

At 10:41 AM, store 3500 aborted due to a sector D0 vacuum problem. Operators called in an expert who discovered a bad vacuum gauge and a vacuum pump off.

Operations put the TeV into shot setup at 2:46 PM.

At 5:12 PM, Operations established store 3504 with an initial luminosity of 48.96E30.

Operations reported at 6 PM that Pbar was off. Operators called in an expert who replaced the phase imbalance circuitry for a power supply (D:H807B). Stacking resumed at 7:25 PM.

A Recycler expert began a study period at 7:46 PM and ended it at 11 PM.

Monday May 17
The midnight shift began with Operations monitoring store 3504, with Pbar stacking, and with MiniBooNE and MCenter taking beam.

At 4 AM, a TeV beam problem appeared. Operators tried adjusting the collimators, but the problem remained. CDF and the D-zero (D0) experiments turned off their detector high voltages. A TeV expert suggested readjusting collimators and that helped a bit, but not enough for the experiment to resume taking data.

A Recycler expert began a study period at 4:34 AM.

Plans for Today
The plans for today are to fix the colliding beam problems, conduct end-of-store and stacking studies later this morning, allow a four to eight hour access into CDF, D0, and the TeV, and then establish a new store.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 5/17/2004   email Fermilab